OpenStudy (anonymous):

The tip of an 11-inch wiper blade wipes a path that is 31 inches long. What is the angle of rotation of the blade in radians to the nearest tenth?

jimthompson5910 (jim_thompson5910):

What is the circumference of this circle (radius = 11)?

OpenStudy (anonymous):

C=2(pi)r c=2(11)pi c-22pi

jimthompson5910 (jim_thompson5910):

22pi, yes

jimthompson5910 (jim_thompson5910):

so if the wiper traced out a path that was exactly 22pi inches, then it would cover the entire circle however, it's just 31 inches, so it covers just a fraction or portion of this circumference

jimthompson5910 (jim_thompson5910):

In radians, a full rotation is 2pi radians we multiply the fraction 31/(22pi) by 2pi to get \[\Large \frac{31}{22\pi}*2\pi = \frac{62\pi}{22\pi} = \frac{31}{11}\]

jimthompson5910 (jim_thompson5910):

So that's the radian measure of the angle. Notice how if we did a full rotation we'd multiply 2pi by (22pi)/(22pi) =1 to get 2pi which is what we should expect

OpenStudy (anonymous):

thank you so much :)